OpenStudy (emilyjones284):

For medal and fan please help (: Write an equation of a parabola that opens upward, has a vertex at the origin, and a focus at (0, 1).

OpenStudy (emilyjones284):

the possible answers are \[y=-\frac{ 1 }{ 4 }x^2, y=\frac{ 1 }{ 4 }x^2, y=\frac{ 1 }{ 12}x^2, y=-\frac{ 1 }{ 12}x^2\]

OpenStudy (whpalmer4):

Vertex form for a parabola is \[y = a(x-h)^2 + k\]where the vertex is at \((h,k)\) That means our equation is of the form \[y = a(x-0)^2 +0\]or\[y = ax^2\]So far, that doesn't help us at all :-) Parabolas that open upward have a value of \(a>0\). We still have 2 choices left. We can find \(a\) by \[a = \frac{1}{4p}\]where \(p\) is the distance from the focus to the vertex. That distance for this parabola is the distance from (0,1) to (0,0), or 1.\[a = \frac{1}{4*1} = \frac{1}{4}\] Therefore, our answer is \[y = \frac{1}{4}x^2\]